OpenStudy (anonymous):

Explain how you would find the exact value of (3)/(√(8)) (3 over the square root of 8)

OpenStudy (anonymous):

Well first you have to find the sqrt of 8

OpenStudy (anonymous):

And that is 2.7

OpenStudy (anonymous):

Wait i feel i am doing this wrong hold on a second

OpenStudy (anonymous):

nvm i cant think right now sorry

OpenStudy (anonymous):

Good luck though

OpenStudy (anonymous):

I know i am a little late but heres what you do

OpenStudy (anonymous):

\[3/\sqrt{8} * \sqrt{8}/\sqrt{8} \]

OpenStudy (anonymous):

You do this because you want to get the 8 out of the square root sign

OpenStudy (anonymous):

Can you figure it out from there if you multiply across?

OpenStudy (theballer225):

dfewff3ferferwferwerwfvrwefrwfrewfwerfwerferwferw